the length of a rectangle is 5 meters less than 4 times the width. if the perimeter is 150 meters, find the length and the width of the rectangle.
Answer by addingup(3677)

L = 4W-5 correct? Make sure you understand what I did. We are going to use this value for L next:
2L+2W = 150 substitute for L:
2(4W-5)+2W = 150 distribute the 2 over the values inside the parenthesis (Distributive Properties of Multiplication):
8W-10+2W = 150 Add W on left, add 10 on both sides and you get:
10W = 160 Divide both sides by 10
W = 16